The demand function for a certain product is given by 

p = 500 + 1000/(q+1)


where p is the price and q is the number of units demanded. Find the average price as demand ranges from 43 to 98 units. (Round your answer to the nearest cent.) 
Answer #1

Ans. =

Price (P1) when 43 Units Quantity Demanded:

P1:

= 500 + [1000 / (q + 1)]
= 500 + [1000 / (43 + 1)]
= 500 + [1000 / 44]
= 500 + 22.72
= $522.72

Price (P2) when 98 Units Quantity Demanded:

P1:

= 500 + [1000 / (q + 1)]
= 500 + [1000 / (98 + 1)]
= 500 + [1000 / 99]
= 500 + 10.10
= $510.10

Average Price:

= (P1 + P2) / 2
= (522.72 + 510.10) / 2
= 1032.82 / 2
= $516.41